OpenStudy (anonymous):

What is the simplified form of this expression 6sqrt v - sqrt 4v^2- sqrt36v+ sqrt v^2? 3v v sqrt v −v 6v2

OpenStudy (cwrw238):

bring like terms together: 6sqrtv - sqrt36v + sqrtv^2 -sqrt4v^2 = 6sqrtv - 6sqrtv + v - 2v - can you continue?

OpenStudy (anonymous):

no :(

OpenStudy (compassionate):

Is this the equation you want solved : \[6\sqrt{v} - \sqrt{4v^2} - \sqrt{36v} + \sqrt{v^2}\]

OpenStudy (compassionate):

I mean expression. The way you wrote it is very sloppy and I'm having trouble deciphering what you mean.

OpenStudy (compassionate):

\[6\sqrt{v} - \sqrt{4v^2} - \sqrt{36v} + \sqrt{v^2} = -2v+v\]